Home‎ > ‎Crew 161‎ > ‎

0112-DailySummaryReport

Executive Officer Report
Crew XO Adeniyi Adebayo

MDRS Daily Summary Report
Sol 3   1/12/2016

Hiking can be fun on Mars.

Executive Officer Reporting

Mission Status: Research projects are on track and mission is going on smoothly.

Sol Activity Summary: Sol 3 was dedicated to collecting more samples from across the Red Planet. Three crew members went on today’s EVA. The team tested radio signal strength as a factor of distance from the HAB. Our crew biologist now has enough samples of different soils with different grades and properties to work on for her plant studies. Crew Astronomer spent the day calibrating the telescope and hopefully, very soon, we will have some beautiful shots from the Martian skies.

Look Ahead Plan: An EVA is planned for tomorrow. Crew commander hopes to explore some further parts east of the HAB. Physical and chemical analysis of the different soils will follow.

Anomalies in work: There are no anomalies to report.

Weather: The Martian sky was very bright and clear today.

Crew Physical Status: All crew members are hale and hearty! We had a routine check-up with our HSO who is satisfied with the physical status of all crew members.

EVA: Three soil samples were collected from across Mars during today’s EVA. ATV locations were noted and marked for future reference.

Reports to be filed:
Commanders Report
EVA #2 report
Daily Summary Report
Astronomy report
Journalist Report
Biologist Report
Engineering Report

Support Requested: Approval for EVA #3

Best Regards,
Adeniyi
Crew Executive Officer